cos(x) / sec(x) + tan(x)
1
Expert's answer
2010-10-11T02:09:15-0400
cos(x) / sec(x) + tan(x) = cos (x) /(1/cos(x)) + sin(x) / cos (x) = cos2(x) + tan(x)